OpenStudy (anonymous):

The slope of the line below is -2. Write the equation of the line in point-slope form, using the coordinates (-2,1)

OpenStudy (anonymous):

Ok so: y-y1=m(x-x1) Tell me what M is, x1 is, & y1 :)

OpenStudy (anonymous):

m is -2 x is 1 and y is -2

OpenStudy (anonymous):

u got the slope right and i think u looked at the point the wrong way again :) it's x=-2 & y=1, remember a point is -> (x1,y1)

OpenStudy (anonymous):

so i got y-1=-2(x-2)

OpenStudy (anonymous):

x1=-2 & y1=1*

OpenStudy (anonymous):

so y+1=-2(x-2)?

OpenStudy (anonymous):

oh whoops sry slow internet, didn't see ur post and let me see

OpenStudy (anonymous):

its all good man

OpenStudy (anonymous):

sort of, just one mistake ok I'll draw it quickly

OpenStudy (anonymous):

alright

OpenStudy (anonymous):